OpenStudy (anonymous):

If the sum of the first two consecutive integers is 120 whay is it

OpenStudy (cj49):

let the consecutive integers be n,n+1 and n+2 n+n+1=4(n+2) 2n+1=4n+8 2n-4n=8-1 -2n=7 n=-7/2

OpenStudy (cj49):

-7/2+(-7/2+1)+(-7/2+2)=?

OpenStudy (cj49):

correct..

OpenStudy (cj49):

the sum of the three intergers is -7.5 and the three integers are -7/2,-5/2 and -3/2

OpenStudy (cj49):

np..